OpenStudy (anonymous):

In triangle DEF, DF = 22 and (angle) F = 45 (degrees). Find the length of a leg. Leave your answer in simplest radical form.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[=\frac{ 22 }{ \sqrt{2} }\] \[=>\frac{ 22 }{ \sqrt{2} }\times\frac{ \sqrt{2} }{ \sqrt{2} }\] \[=>\frac{ 22\sqrt{2} }{ 2 }\] \[=>12\sqrt{2}\]

OpenStudy (anonymous):

Don't you divide 22 by 2? @r4fay

OpenStudy (anonymous):

I made a mistake there. It should be 11sqrt(2) instead of 12sqrt(2).

OpenStudy (anonymous):

Ah I see. It's okay. Thank you so much! @r4fay